OpenStudy (anonymous):

will give medal for explaination A computer downloads files at a constant rate. The table shows how many megabytes the computer downloads over specific lengths of time. What is the slope of the line that represents the computer downloads? Computer Downloads Minutes 3 5 7 9 Megabytes 11.25 18.75 26.25 33.75 i think its 3.75

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

\[Slope = \frac{\textsf{Change \in Megabytes}}{\textsf{Change \in Minutes}}\]

OpenStudy (anonymous):

\[Slope = \frac{18.75 - 11.25}{5-3} \implies Slope = \frac{7.5}{2} = ??\]

OpenStudy (anonymous):

Yeah, that is looking fine to me.. Good..!! :)

OpenStudy (anonymous):

so its 3.75?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Have confidence on yourself.. :) I said, \(\huge YES\).. :)

OpenStudy (anonymous):

ok thank you